The market requires a return of 9% from XYZ, Inc. The firm plowback 50% of its earnings, and its return on equity and earnings per share are expected to be 14% and $7, respectively.

a. What will be XYZ's growth rate? (Input your answer as a nearest whole percent.)


b. Calculate XYZ's P/E ratio? (Do not round intermediate calculations.)
Answer #1

a.

Growth Rate = Plowback ratio * Return on Equity

= 0.50 * 0.14

Growth Rate = 0.07 or 7%

b.

Plowback Ratio = 1 - (DPS/EPS)

0.50 = 1 - (DPS/7)

DPS = (1 - 0.50) * 7

DPS = 3.5

Market Price = D0(1+g) / (k - g)

= 3.5(1+0.07) / (0.09 - 0.07)

Market Price = $187.25

P/E Ratio = Price per share / Earning per share

= 187.25 / 7

P/E Ratio = 26.75